OpenStudy (anonymous):

How to do substitution in algebra like 5c-d-11=0,4d+3c=-5

OpenStudy (anonymous):

Can any one help ? I have a exam tomorrow

OpenStudy (shinalcantara):

since it's given that you have 2 variables which doesn't have values, then substitution is a good idea of solving for the values. 5c-d-11 = 0 -------eqn 1 4d + 3c =-5 -------eqn 2 first, try to examine which of the two is simple such that when you'll find it's value in terms of the other variable you'll have a simple equation. Let's choose eqn 1 since d has 1 as a coefficient 5c-d-11=0 now transpose d to the other side making it 5c-11=d or d = 5c-11 -----eqn 3 now you are ready to substitute eqn 3 to eqn 2 4d + 3c = -5 since you have the value for d from eqn 3, you'll have 4(5c-11) + 3c = -5 now simplify that equation and you'll get the value for c

OpenStudy (shinalcantara):

once you'll have the value for c, you can find d by substituting c's value to eqn 3

OpenStudy (anonymous):

Thanks :)

OpenStudy (shinalcantara):

yw :)